# Decentralized Proactive Model Offloading and Resource Allocation for Split and Federated Learning

### v1

@@ 7.2 Performance evaluation of the proposed DP-MORA scheme
1) Training latency: Fig. 2 shows the per-round training latency of different schemes over Resnet 18 and Resnet 34 two models when the data leakage risk rate constraint $P^{risk}$ is 0.5. We can see from Fig. 2 that the per-round training latency of the DP-MORA scheme is $24.95\%$ lower than that of SF3AF, $24.09\%$ lower than that of FAAF, $31.72\%$ lower than that of SF3PF and FSPF, $86.02\%$ lower than that of SF1AF, $86.35\%$ lower than that SF1PF, $84.56\%$ lower than that of SF2AF, $85.14\%$ lower than that SF2PF, and $24.09\%$ lower than that of FSAF when the data leakage risk rate is less than or equal to 0.5. The per-round training latency consists of device-side model distribution latency, training latency of five epochs and device-side model transmission latency. That is because that considering the heterogeneity of IoT devices’ resource capacities and their local data samples’ size, the DP-MORA scheme solves the optimal model offloading and resource allocation scheme which make the standard deviation of per-epoch training latencies of IoT devices as small as possible. Therefore, the DP-MORA scheme can mitigate the straggler effect and reduce the per-epoch training latency, thereby reducing the per-round training latency.
@@
The per-round training latency of SF1AF and SF1PF are the highest. There are three main reasons: (1) SF1AF and SF1PF split the DNN model on each IoT device at the same layer without considering the heterogeneity of IoT devices’ resource capacities and their local data samples’ size; (2) SF1AF and SF1PF collaborate multiple IoT devices with the edge server to train DNN model in a sequential way, which incur long training latency; (3) SF1AF and SF1PF allocate the shared network resource and computing resource of the edge server in average and proportional way without considering the impact of efficient resource allocation on the training latency.
@@
The per-round training latency of SF2AF and SF2PF are the second highest. There are two main reasons: (1) although they adopt the same model offloading strategy as the DP-MORA scheme, SF2AF and SF2PF collaborate multiple IoT devices with the edge server to train DNN model in a sequential way; (2) SF2AF and SF2PF allocate the shared network resource and computing resource of the edge server in average and proportional way without considering the impact of efficient resource allocation on the training latency.
@@
The per-round training latency of SF3AF and SF3PF are higher than that of the DP-MORA scheme. The main reason is that although SF3AF and SF3PF adopt the same model offloading strategy as the DP-MORA scheme, they do not consider the heterogeneity of IoT devices’ resource capacities and their local data samples’ size, but also optimal allocation of bandwidth and computing resources.
@@
FSAF and FSPF offload as many layers as possible to the edge server while satisfying the data leakage risk rate constraint 0.5. However, the per-round training latency of FSAF and FSPF are higher that of the DP-MORA scheme. That is because that FSAF and FSPF split each IoT device’s DNN model at the same cut layer without considering the heterogeneity of IoT devices’ resource capacities and their local data samples’ size. In addition, FSAF and FSPF also do not consider the impact of edge server’s resource allocation on the training latency.
@@
FAAF locally trains the whole DNN model on resource-constrained IoT devices. Therefore, its data leakage risk rate is lowest. However, the per-round training latency of FAAF is higher than that of the DP-MORA scheme. That is because that FAAF does not fully exploit the resource-adequate edge server to achieve device-edge synergy training, thereby incurring longer per-round training latency.
@@
2) Waiting latency: After multiple epochs training of each IoT device, its latest device-side model is transmitted to the edge server to further perform model aggregation. As a result, the earliest IoT device with its device-side model transmitted to the edge server needs to wait for the latest IoT device. For the sake of convenience, the waiting latency of an IoT device is defined as the duration between its device-side model uploaded to the edge server and the last IoT device’s model uploaded to the edge server. Table. II and Table. III illustrates each IoT device’s waiting latency of different schemes over Resnet 18 and Resnet 34 two models when the data leakage risk rate constraint $P^{risk}$ is 0.5. We can observe from Table. II and Table. III that the IoT devices’ waiting latencies in DP-MORA fluctuate slightly. That is because that the DP-MORA scheme can identify the optimal model offloading and resource allocation strategy which allocate the powerful IoT device with less local data samples less bandwidth resources and computation resources, and allocate the less capable IoT device with more local data samples more bandwidth resources and computation resources, thereby thereby reducing the waiting latency for less capable IoT device with more local data samples and improving the training efficiency. Otherwise, if the bandwidth resource and computing resource of the edge server are equally divided to each IoT device, the weaker IoT device with more local data samples need to spend more time in training models, inevitably incurring more waiting latency and the total training latency, thereby reducing the training efficiency.
@@
We can also see from Table. II and Table. III that the waiting latency of the DP-MORA scheme is lowest, the waiting latency of SF3AF, FSAF and FAAF are medium, the waiting latency of SF1AF and SF2AF are longest. The reason for the DP-MORA scheme with lowest waiting latency is that the DP-MORA scheme can identify the optimal model offloading and resource allocation scheme considering the heterogeneity of IoT devices’ resource capacities and their local data samples’ size. The reason for FSAF and FAAF with the highest waiting latency is that they collaborate multiple IoT devices with the edge server to train DNN model in a sequential way, thereby incurring long training latency. The reason for SF3AF, FSAF and FAAF with the medium waiting latency is that they train DNN model in a parallel manner. Specifically, SF3AF adopt the same model offloading strategy as the DP-MORA, but it does not consider the impact of efficient resource allocation on training latency. FSAF offloads the identical server-side sub-model to the edge server for device-edge synergy training in a parallel manner. However, it not only fails to consider the heterogeneity of IoT devices’ resource capacities and their local data samples’ size, but also fails to consider the edge server’s resource allocation. FAAF train the whole DNN model on local IoT devices in parallel. However, it does not fully exploit the resource-adequate edge server to achieve efficient device-edge synergy training.
TABLE II: Per-round waiting latency of each IoT device over Resnet 18 model
IoT Devices Approaches
@@
TABLE III: Per-round per-waiting latency of each IoT device over Resnet 34 model
IoT Devices Approaches
@@
3) Model accuracy: To verify the impact of the proposed DP-MORA scheme on the model accuracy, we plot its model accuracy curves over Resnet 18 and Resnet 34 with the increase of the training rounds and the training latency. Fig. 3 and Fig. 4 illustrates the model accuracy of different schemes over Resnet 18 and Resnet 34 two models, when the data leakage risk rate constraint $P^{risk}$ is 0.5. We can observe from Fig. 3 and Fig. 4 that the model accuracy of the DP-MORA scheme gradually increases and eventually converges with the increase of the training rounds. In addition, we can observe that the model accuracy of the DP-MORA scheme is consistent with that of FAAF. The main reason is that the DP-MORA scheme mainly optimize on training efficiency by jointly model offloading and resource allocation, which has no influence on model accuracy.
@@
Figure 3: Accuracies of different approaches over Restnet 18 on cifar10 dataset
Figure 4: Accuracies of different approaches over Restnet 34 on cifar10 dataset
4) Impact of different data leakage risk rate constraints: To examine the impact of different data leakage risk rate constraints on the total training latency, we vary the data leakage risk rate constraint from 0.1 to 0.8 with the increment of 0.1. Fig. 5 plots the related experimental results. In Fig. 5, we can see that the total training latency of the DP-MORA scheme gradually decreases as the data leakage risk rate constraint increases. That is because different data leakage risk rate constraints correspond to different cut layers requirement. The higher the data leakage risk rate constraint is, the shallower the cut layer is, thereby the solution space of the optimization problem with minimizing training latency while satisfying the data leakage risk rate constraint is larger. The DP-MORA scheme has a higher probability to identify the optimal solution from the larger feasible solution space. The optimal solution is the optimal cut layer and resource requirement for each IoT device. The optimal solution splits the DNN model at an optimal shallow cut layer and offload the server-side sub-model with more layers to the resource-adequate edge server, and allocates optimal bandwidth resource for the transmission data and optimal computation resource for the server-side sub-model, thereby obtaining lower training latency. On the contrary, when the data leakage risk rate constraint is low, it indicates that IoT devices have a low tolerance for data leakage. Based on this, only several deep cut layers of the DNN model can be selected to meet the data leakage risk rate constraint. A deep cut layer means heavy computation workloads on resource-constrained IoT devices and light computational workload on resource-adequate edge server, which incurring a longer training latency.
Moreover, we can further observe that compared to SF3AF, SF3PF and FAAF benchmarks, the DP-MORA scheme achieves the lowest total training latency while satisfying the data leakage risk rate constraints. Specifically, when data leakage risk rate constraint is 0.8, compared to SF3AF, SF3PF and FAAF, the DP-MORA scheme reduces the total training latency of Resnet18 model up to $24.95\%$, $31.72\%$ and $24.09\%$, respectively. That is because that the DP-MORA scheme identify the optimal model offloading and resource allocation strategy for heterogeneous IoT devices to reduce the total training latency.
@@
5) Impact of edge server’s different computation capacities: We investigate the impact of edge server’s different computation capacities on the training latency. We vary the computing capacity of the edge server from 50GFLOPS to 150GFLOPS with the increment of 50. The related experimental result is given in Fig. 6. We can see from Fig. 6 that when the data leakage risk rate constraint $P^{risk}$ is 0.5, the per-round training latency of DP-MORA, SF3AF, FSAF, SF2AF and SF1AF schemes over Resnet 18 and Restnet 34 two models gradually decrease with the increase of the edge server’s computation capacity. The main reason is that the edge server with the higher computation capacity can cooperate IoT devices to handle more workloads, and thereby reducing the per-round training latency. The per-round training latency of FAAF over Resnet 18 and Restnet 34 two models are constant with the increase of the edge server’s computation capacity. That is because that FAAF train the whole DNN model on local IoT devices in parallel. Therefore, the increase of the edge server’s computation capacity has no effect on the training latency of FAAF.
We can further observe from Fig. 6 that the training latency of the DP-MORA scheme is lower than those of other baseline algorithms. Specifically, when the computation capacities is 150GFLOPS, the per-round training latency of the DP-MORA scheme is $31.5\%$ lower than that of FAAF, $21.1\%$ lower than that of FSAF and SF3AF, $84.1\%$ and $85.3\%$ lower than that of SF2AF and SF1AF when the data leakage risk rate constraint 0.5 is satisfied on Resnet18 model. And when the computation capacities is 50GFLOPS, the per-round training latency of the DP-MORA scheme is $26\%$ lower than that of FAAF, $33.1\%$ lower than that of SF3AF and FSAF, $85.4\%$ lower than that of SF2AF, $87.9\%$ lower than that of SF1AF when the data leakage risk rate constraint 0.5 is satisfied on Resnet18 model. That is because that the edge server’s computation resource becomes more and more adequate with the increase of the edge server’s computation capacity. The DP-MORA scheme enable to take full usage of more adequate computation resource to collaborate multiple heterogeneous IoT devices to efficiently train DNN model, thereby greatly reducing the training latency.
@@
6) Impact of edge server’s uplink bandwidth: We evaluate the impact of edge server’s uplink bandwidth on the training latency. We vary the edge server’s uplink bandwidth from 100Mbps to 400Mbps. The related experimental result is shown in Fig. 7. We see that the per-round training latency of the DP-MORA scheme and other baseline algorithms gradually decrease. That is because that when the edge server’s uplink bandwidth increases, the time taken to transmit the same amount of data can decrease. Moreover, when the edge server’s uplink bandwidth increases, we can also see that the per-round training latency of the proposed DP-MORA scheme is always lower than that of other benchmarks. Specifically, when the edge server’s uplink bandwidth is 100Mbps, the DP-MORA scheme reduces the per-round training latency by $36.6\%$, $27.4\%$, and $27.4\%$ as compared with the FAAF, SF3AF, and FSAF benchmarks, respectively. The DP-MORA scheme is $85\%$ lower than that of SF2AF, $86.7\%$ lower than that of SF1AF due to the sequential execution instinct for SF1AF and SF2AF. When the edge server’s uplink bandwidth is 400Mbps, the DP-MORA scheme reduces the per-round training latency by $31.2\%$, $22.5\%$, and $22.5\%$ as compared with the FAAF, SF3AF, and FSAF benchmarks, respectively. DP-MORA is $84.5\%$ lower than that of SF2AF, $85.6\%$ lower than that of SF1AF. That is because that the DP-MORA scheme optimally allocate edge server’s uplink bandwidth resources for multiple heterogeneous IoT devices, thereby reducing the training latency.
@@
7) Impact of edge server’s downlink bandwidth: In order to examine the impact of edge server’s downlink bandwidth on the per-round training latency, we vary the edge server’s downlink bandwidth from 50Mbps to 200Mbps. The related experimental results are shown in Fig. 8. We see that when the edge server’s downlink bandwidth increases, the per-round training latency of the DP-MORA scheme and other benchmarks decreases. The main reason is that it takes less time to transmit the same amount of data with the larger uplink bandwidth. Moreover, we can further see from Fig. 8 that the per-round training latency of the DP-MORA scheme is lower than that of other benchmarks. Specifically, when the edge server’s downlink bandwidth is 50Mbps on Resnet18 model, the DP-MORA is $25.6\%$ lower than that of FAAF, $31.4\%$ lower than that of SF3AF and FSAF, $85.1\%$ lower than that of SF2AF, $87.6\%$ lower than that of SF1AF. When the edge server’s downlink bandwidth is 200Mbps, the DP-MORA is $29.9\%$ lower than that of FAAF, $17.5\%$ lower than that of SF3AF and FSAF, $83.6\%$ lower than that of SF2AF, $84.5\%$ lower than that of SF1AF. That is because that considering the heterogeneity of IoT devices’ resource capacities and their local data samples’ size, the DP-MORA scheme can optimally allocate the downlink bandwidth resource to each IoT device.
@@
8 Conclusion
In this paper, we investigate the training efficiency of split federated (SplitFed) learning while satisfying the data leakage risk rate constraint in resource-constrained IoT-edge environment. To address this problem, we first formulate the latency of DNN model training and measure the data leakage risk rate of DNN model training adopting split federated learning. Then we formulate joint model offloading and resource allocation problem to be a mixed integers non-linear programming. At last, we design a decentralized and privacy-preserving joint model offloading and resource allocation scheme to optimize the per-round training latency while satisfying data leakage risk rate constraint. We conduct extensive experiments on two real-world datasets to evaluate the performance. Extensive experiments show that the DP-MORA scheme can effectively reduce the per-round training latency while satisfying the data leakage risk rate constraint.

### v2

@@ VII-B Performance evaluation of the proposed DP-MORA scheme
1) Training latency: Fig. 2 shows the per-round training latency of different schemes over ResNet 18 and ResNet 34 two models when the data leakage risk rate constraint $P^{risk}$ is 0.5. We can see from Fig. 2 that the per-round training latency of the DP-MORA scheme is $24.95\%$ lower than that of SF3AF, $24.09\%$ lower than that of FAAF, $31.72\%$ lower than that of SF3PF and FSPF, $86.02\%$ lower than that of SF1AF, $86.35\%$ lower than that SF1PF, $84.56\%$ lower than that of SF2AF, $85.14\%$ lower than that SF2PF, and $24.09\%$ lower than that of FSAF when the data leakage risk rate is less than or equal to 0.5. The per-round training latency consists of device-side model distribution latency, training latency of five epochs and device-side model transmission latency. That is because considering the heterogeneity of end devices’ resource capacities and their local data samples’ size, the DP-MORA scheme solves the optimal model offloading and resource allocation scheme which makes the standard deviation of per-epoch training latencies of end devices as small as possible. Therefore, the DP-MORA scheme can mitigate the straggler effect and reduce the per-epoch training latency, thereby reducing the per-round training latency.
@@
The per-round training latency of SF1AF and SF1PF is the highest. There are three main reasons: (1) SF1AF and SF1PF split the DNN model on each end device at the same layer without considering the heterogeneity of end devices’ resource capacities and their local data samples’ size; (2) SF1AF and SF1PF collaborate multiple end devices with the edge server to train DNN model in a sequential way, which incur high training latency; (3) SF1AF and SF1PF allocate the shared network resource and computing resource of the edge server in average and proportional way without considering the impact of efficient resource allocation on the training latency.
@@
The per-round training latency of SF2AF and SF2PF is the second highest. There are two main reasons: (1) although they adopt the same model offloading strategy as the DP-MORA scheme, SF2AF and SF2PF collaborate multiple end devices with the edge server to train DNN model in a sequential way; (2) SF2AF and SF2PF allocate the shared network resource and computing resource of the edge server in an average and proportional way without considering the impact of efficient resource allocation on the training latency.
@@
The per-round training latency of SF3AF and SF3PF is higher than that of the DP-MORA scheme. The main reason is that although SF3AF and SF3PF adopt the same model offloading strategy as the DP-MORA scheme, they do not consider the heterogeneity of end devices’ resource capacities and their local data samples’ size, but also optimal allocation of bandwidth and computing resources.
@@
FSAF and FSPF offload as many layers as possible to the edge server while satisfying the data leakage risk rate constraint 0.5. However, the per-round training latency of FSAF and FSPF is higher that of the DP-MORA scheme. That is because FSAF and FSPF split each end device’s DNN model at the same cut layer without considering the heterogeneity of end devices’ resource capacities and their local data samples’ size. In addition, FSAF and FSPF also do not consider the impact of the edge server’s resource allocation on the training latency.
@@
FAAF locally trains the whole DNN model on resource-constrained end devices. Therefore, its data leakage risk rate is the lowest. However, the per-round training latency of FAAF is higher than that of the DP-MORA scheme. That is because FAAF does not fully exploit the resource-adequate edge server to achieve device-edge synergy training, thereby incurring higher per-round training latency.
@@
2) Waiting latency: After multiple epochs of training of each end device, its latest device-side model is transmitted to the edge server to further perform model aggregation. As a result, the earliest end device with its device-side model transmitted to the edge server needs to wait for the latest end device. For the sake of convenience, the waiting latency of an end device is defined as the duration between its device-side model uploaded to the edge server and the last end device’s model uploaded to the edge server. Table. III and Table. IV illustrate each end device’s waiting latency of different schemes over ResNet 18 and ResNet 34 two models when the data leakage risk rate constraint $P^{risk}$ is 0.5. Based on Table. III, the per-round waiting latency variances of SF1AF, DP-MORA, SF2AF, SF3AF, FSAF and FAAF six schemes over ResNet 18 are 6427.60, 78.18 5211.61, 171.48, 123.32, 211.14, respectively. Based on Table. IV, the per-round waiting latency variances of SF1AF, DP-MORA, SF2AF, SF3AF, FSAF and FAAF six schemes over ResNet 18 are 17568.76, 192.118189, 16475.28, 380.02, 326.48, 856.96, respectively. Therefore, the end devices’ waiting latencies in DP-MORA fluctuate slightly. That is because the DP-MORA scheme can identify the optimal model offloading and resource allocation strategy. This strategy allocates less bandwidth resources and computing resources of edge server to powerful end devices with fewer samples. Conversely, it allocates more bandwidth resources and computing resources of edge server to those weak end devices with more samples, thereby reducing the waiting latency for less capable end devices with more local data samples and improving the training efficiency. Otherwise, if the bandwidth resource and computing resource of the edge server are equally divided to each end device, the weaker end device with more local data samples needs to spend more time in training models, inevitably incurring more waiting latency and the total training latency, thereby reducing the training efficiency.
@@
We can also see from Table. III and Table. IV that the waiting latency of the DP-MORA scheme is lowest, the waiting latency of SF3AF, FSAF and FAAF are medium, the waiting latency of SF1AF and SF2AF is highest. The reason for the DP-MORA scheme with lowest waiting latency is that the DP-MORA scheme can identify the optimal model offloading and resource allocation scheme considering the heterogeneity of end devices’ resource capacities and their local data samples’ size. The reason for FSAF and FAAF with the highest waiting latency is that they collaborate multiple end devices with the edge server to train DNN model in a sequential way, thereby incurring high training latency. The reason for SF3AF, FSAF and FAAF with the medium waiting latency is that they train DNN model in a parallel manner. Specifically, SF3AF adopts the same model offloading strategy as the DP-MORA scheme, but it does not consider the impact of efficient resource allocation on training latency. FSAF offloads the identical server-side sub-model to the edge server for device-edge synergy training in a parallel manner. However, it not only fails to consider the heterogeneity of end devices’ resource capacities and their local data samples’ size but also fails to consider the edge server’s resource allocation. FAAF trains the whole DNN model on local end devices in parallel. However, it does not fully exploit the resource-adequate edge server to achieve efficient device-edge synergy training.
TABLE III: Per-round waiting latency of each end device over
ResNet 18 model
End Devices Approaches
@@
TABLE IV: Per-round waiting latency of each end device over
ResNet 34 model
End Devices Approaches
@@
3) Model accuracy: To verify the impact of the proposed DP-MORA scheme on the model accuracy, we plot its model accuracy curves over ResNet 18 and ResNet 34 with the increase of the training rounds and the training latency. Fig. 3 and Fig. 4 illustrate the model accuracy of different schemes over ResNet 18 and ResNet 34 two models, when the data leakage risk rate constraint $P^{risk}$ is 0.5. We can observe from Fig. 3 and Fig. 4 that the model accuracy of the DP-MORA scheme gradually increases and eventually converges with the increase of the training rounds. In addition, we can observe that the model accuracy of the DP-MORA scheme is consistent with that of FAAF. The main reason is that the DP-MORA scheme mainly optimizes on training efficiency by jointly model offloading and resource allocation, which has no influence on model accuracy.
@@
Figure 3: Accuracies of different approaches over ResNet 18 on cifar10 dataset
Figure 4: Accuracies of different approaches over ResNet 34 on cifar10 dataset
4) Impact of different data leakage risk rate constraints: To examine the impact of different data leakage risk rate constraints on the total training latency, we vary the data leakage risk rate constraint from 0.1 to 0.8 with the increment of 0.1. Fig. 5 plots the related experimental results. In Fig. 5, we can see that the total training latency of the DP-MORA scheme gradually decreases as the data leakage risk rate constraint increases. That is because different data leakage risk rate constraints correspond to different cut layers requirement. The higher the data leakage risk rate constraint is, the shallower the cut layer is, thereby the solution space of the optimization problem with minimizing training latency while satisfying the data leakage risk rate constraint is larger. The DP-MORA scheme has a higher probability of identifying the optimal solution from the larger feasible solution space. The optimal solution is the optimal cut layer and resource requirements for each end device. The optimal solution splits the DNN model at an optimal shallow cut layer and offloads the server-side sub-model with more layers to the resource-adequate edge server, and allocates optimal bandwidth resource for the transmission data and optimal computation resource for the server-side sub-model, thereby obtaining lower training latency. On the contrary, when the data leakage risk rate constraint is low, it indicates that end devices have a low tolerance for data leakage. Based on this, only several deep-cut layers of the DNN model can be selected to meet the data leakage risk rate constraint. A deep cut layer means heavy computation workloads on resource-constrained end devices and light computational workload on resource-adequate edge server, which incur a higher training latency.
Moreover, we can further observe that compared to SF3AF, SF3PF and FAAF benchmarks, the DP-MORA scheme achieves the lowest total training latency while satisfying the data leakage risk rate constraints. Specifically, when the data leakage risk rate constraint is 0.8, compared to SF3AF, SF3PF and FAAF, the DP-MORA scheme reduces the total training latency of the ResNet18 model up to $24.95\%$, $31.72\%$ and $24.09\%$, respectively. That is because the DP-MORA scheme identifies the optimal model offloading and resource allocation strategy for heterogeneous end devices to reduce the total training latency.
@@
5) Impact of edge server’s different computation capacities: We investigate the impact of edge server’s different computation capacities on the training latency. We vary the computing capacity of the edge server from 50GFLOPS to 150GFLOPS with the increment of 50. The related experimental result is given in Fig. 6. We can see from Fig. 6 that when the data leakage risk rate constraint $P^{risk}$ is 0.5, the per-round training latency of DP-MORA, SF3AF, FSAF, SF2AF and SF1AF schemes over ResNet 18 and ResNet 34 two models gradually decrease with the increase of the edge server’s computation capacity. The main reason is that the edge server with the higher computation capacity can cooperate end devices to handle more workloads, and thereby reducing the per-round training latency. The per-round training latency of FAAF over ResNet 18 and ResNet 34 two models are constant with the increase of the edge server’s computation capacity. That is because FAAF trains the whole DNN model on local end devices in parallel. Therefore, the increase of the edge server’s computation capacity has no effect on the training latency of FAAF.
We can further observe from Fig. 6 that the training latency of the DP-MORA scheme is lower than those of other baseline algorithms. Specifically, when the computation capacity is 150GFLOPS, the per-round training latency of the DP-MORA scheme is $31.5\%$ lower than that of FAAF, $21.1\%$ lower than that of FSAF and SF3AF, $84.1\%$ and $85.3\%$ lower than that of SF2AF and SF1AF when the data leakage risk rate constraint 0.5 is satisfied on ResNet18 model. And when the computation capacity is 50GFLOPS, the per-round training latency of the DP-MORA scheme is $26\%$ lower than that of FAAF, $33.1\%$ lower than that of SF3AF and FSAF, $85.4\%$ lower than that of SF2AF, $87.9\%$ lower than that of SF1AF when the data leakage risk rate constraint 0.5 is satisfied on ResNet18 model. That is because the edge server’s computation resource becomes more and more adequate with the increase of the edge server’s computation capacity. The DP-MORA scheme enables to take full use of more adequate computation resources to collaborate with multiple heterogeneous end devices to efficiently train DNN model, thereby greatly reducing the training latency.
@@
6) Impact of edge server’s uplink bandwidth: We evaluate the impact of edge server’s uplink bandwidth on the training latency. We vary the edge server’s uplink bandwidth from 100Mbps to 400Mbps. The related experimental result is shown in Fig. 7. We see that the per-round training latency of the DP-MORA scheme and other baseline algorithms gradually decrease. That is because when the edge server’s uplink bandwidth increases, the time taken to transmit the same amount of data can decrease. Moreover, when the edge server’s uplink bandwidth increases, we can also see that the per-round training latency of the proposed DP-MORA scheme is always lower than that of other benchmarks. Specifically, when the edge server’s uplink bandwidth is 100Mbps, the DP-MORA scheme reduces the per-round training latency by $36.6\%$, $27.4\%$, and $27.4\%$ as compared with the FAAF, SF3AF, and FSAF benchmarks, respectively. The DP-MORA scheme is $85\%$ lower than that of SF2AF, and $86.7\%$ lower than that of SF1AF due to the sequential execution instinct for SF1AF and SF2AF. When the edge server’s uplink bandwidth is 400Mbps, the DP-MORA scheme reduces the per-round training latency by $31.2\%$, $22.5\%$, and $22.5\%$ as compared with the FAAF, SF3AF, and FSAF benchmarks, respectively. DP-MORA is $84.5\%$ lower than that of SF2AF, and $85.6\%$ lower than that of SF1AF. That is because the DP-MORA scheme optimally allocates edge server’s uplink bandwidth resources for multiple heterogeneous end devices, thereby reducing the training latency.
@@
7) Impact of edge server’s downlink bandwidth: In order to examine the impact of edge server’s downlink bandwidth on the per-round training latency, we vary the edge server’s downlink bandwidth from 50Mbps to 200Mbps. The related experimental results are shown in Fig. 8. We see that when the edge server’s downlink bandwidth increases, the per-round training latency of the DP-MORA scheme and other benchmarks decreases. The main reason is that it takes less time to transmit the same amount of data with the larger uplink bandwidth. Moreover, we can further see from Fig. 8 that the per-round training latency of the DP-MORA scheme is lower than that of other benchmarks. Specifically, when the edge server’s downlink bandwidth is 50Mbps on the ResNet18 model, the DP-MORA is $25.6\%$ lower than that of FAAF, $31.4\%$ lower than that of SF3AF and FSAF, $85.1\%$ lower than that of SF2AF, $87.6\%$ lower than that of SF1AF. When the edge server’s downlink bandwidth is 200Mbps, the DP-MORA is $29.9\%$ lower than that of FAAF, $17.5\%$ lower than that of SF3AF and FSAF, $83.6\%$ lower than that of SF2AF, $84.5\%$ lower than that of SF1AF. That is because considering the heterogeneity of end devices’ resource capacities and their local data samples’ size, the DP-MORA scheme can optimally allocate the downlink bandwidth resource to each end device.
@@
VIII Conclusion
In this paper, we investigate the training efficiency of split federated (SplitFed) learning while satisfying the data leakage risk rate constraint in resource-constrained IoT-edge computing environment. To address this problem, we first formulate the latency of DNN model training and measure the data leakage risk rate of DNN model training adopting split federated learning. Then we formulate joint model offloading and resource allocation problem to be a mixed integers non-linear programming. At last, we design a decentralized and privacy-preserving joint model offloading and resource allocation scheme to optimize the per-round training latency while satisfying data leakage risk rate constraint. We conduct extensive experiments on two real-world datasets to evaluate the performance. Extensive experiments show that the DP-MORA scheme can effectively reduce the per-round training latency while satisfying the data leakage risk rate constraint.